# Dynamic Decoupling in Multidimensional Screening

### v1

@@ 3 Dynamic (Ir)relevance
st. $\gamma\in\operatorname*{\arg\!\max}\mathbb{E}_{\theta|\gamma}\left[u(\gamma^{\prime},\theta)\right]-t_{1}(\gamma^{\prime})$ $IC1$
@@
where $u(\gamma,\theta)=\theta\cdot q(\gamma,\theta)-t_{2}(\gamma,\theta)$ is the agent’s utility net of $t_{1}$ and $\nabla_{\theta}u(\gamma,\theta)$ is the gradient of $u(\gamma,\theta)$ in $\theta$ at some fixed $\gamma$.
@@
$IC1$ continues to pin down incentive compatibility for truthful reporting of $\gamma$ while $IC2$ captures incentive compatibility for truthful reporting of $\theta$. Next, $F$ is feasibility: The gradient of the agent’s post-contractual utility function $u$ is equal to the allocation rule which must lie between zero and one. Finally, $IR$ pins down interim individual rationality to accept the contract. Total surplus generated by the monopolist is $\theta\cdot\nabla_{\theta}u(\gamma,\theta)$ while a share $u(\gamma,\theta)\frac{f_{\gamma}(\theta|\gamma)}{f(\theta|\gamma)}\frac{1-G(\gamma)}{g(\gamma)}$ is paid to the agent as information rents. The Lemma establishes that the interim utility function $u$ pins down both period one and period two transfers and specifies how dynamics distorts designer profits.
@@
Compared to the usual multi-good monopolist problem where the agent receives their full information rents $u$, allowing the principal to frontload surplus extraction leads to period two information rents given to the agent being extracted by period one transfers. [^3] In the proof of the Lemma, we explicitly work out how the cancellation happens. As such, the only information rents the principal needs to pay the agent are those to truthfully elicit their pre-contractual private information. These information rents are distorted by two factors. First, $\frac{f_{\gamma}(\theta|\gamma)}{f(\theta|\gamma)}=\frac{\partial}{\partial\gamma}\ln(f(\theta|\gamma))=s(\gamma,\theta)$ is the score function measuring how much the log-likelihood of $\theta$ changes as $\gamma$ changes. This captures how likely $\theta$ came from one value of $\gamma$ versus another and thus governs how post-contractual utilities are passed through into information rents. The other $\frac{1-G(\gamma)}{g(\gamma)}$ term is the standard hazard rate; as usual, providing information rents to lower types once again leads to higher information rents to higher types.
@@
Applying the divergence theorem as in opttransport allows us to express $\nabla_{\theta}u(\gamma,\theta)$ in terms of $u(\gamma,\theta)$. However, this significantly complicates the expression and does not offer much additional insight; in fact, it may be simpler to express the problem in terms of $\nabla_{\theta}u(\gamma,\theta)$ instead. There are two further complications preventing us from using an optimal transport or duality (see kleiner2019) approach to analyze the problem. First, as segal2014 notes, the consumer’s expected interim utility function may be discontinuous in $\gamma$. Unlike continuity in $\theta$ being guaranteed by feasibility of the allocation rule, the principal may choose to give information rents to consumers with different values of $\gamma$ at different values of $\theta$. While is is possible to use an orthogonalization approach to transform the problem, making $\mathbb{E}_{\theta|\gamma}\left[u(\gamma,\theta)\right]$ continuous in $\gamma$, it is unclear if $u(\gamma,\theta)$ itself can be made continuous in $\gamma$; or if doing so would be an useful exercise. Second, the usual single-crossing conditions that lead to monotonicity being a necessary and sufficient condition for implementability do not hold. We can view the problem as the principal designing a utility function $w$ where an agent with private information $\gamma$ values $w$ at
$\hat{U}(\gamma,w(\theta))=\int_{\Theta}w(\theta)dF(\theta|\gamma).$
Then, the usual characterizations of single crossing or monotonicity in $w$ cannot be applied as $w$ is a high-dimensional object. However, cyclic monotonicity as in rochet1987 is still a necessary and sufficient condition for implementability, but is difficult to operationalize.
When does Lemma 1 help? Consider the case when different dimensions of $\theta$ are independent (the following intuition generalizes): Then, multidimensional sequential screening problems have a simple solution. In this case, additivity of scores $s(\gamma,\theta)$ leads to independently selling each good being optimal: Each dimension of $\theta$ provides information about $\gamma$ separately. The joint forces of intertemporal screening and independence leads to the solution being separable: Intertemporal screening enables the principal to push surplus extraction to the first period while additivity of scores enables information rents to be computed good-by-good.
@@
$\nabla_{\theta}\cdot(\mathbf{V}(\theta|\gamma)f(\theta|\gamma))=f_{\gamma}(\theta|\gamma).$ (1)
@@
In the independent case, $\mathbf{V}(\theta|\gamma)$ has a simple form: The $j$th component of $\mathbf{V}$ is
$\mathbf{V}^{j}(\theta|\gamma)=\frac{F^{j}_{\gamma}(\theta^{j}|\gamma)}{f^{j}(\theta^{j}|\gamma)}.$ (3)
This solution to $\mathbf{V}$ works in more general settings than perfect conditional independence. By Sklar’s Theorem, there exists an unique distribution $C:[0,1]^{n}\times\Gamma\to[0,1]$ such that
and letting $c$ be the density of $C$,
Definition 1 (Invariant Dependencies).
Distribution $F(\theta|\gamma)$ has invariant dependencies if $c(\cdot|\gamma)$ is constant in $\gamma$. If this is the case, we write $C(\cdot|\gamma)=C(\cdot),c(\cdot|\gamma)=c(\cdot)$.
If the dimensions of $F$ are independent conditional on $\gamma$, then $C$ is uniform over $0,1]^{n}$ for each $\gamma$ so $F$ has invariant dependencies. As such, this condition is a generalization of conditional independence; any choice of marginals $F^{1}(\theta^{1}|\gamma),...,F^{n}(\theta^{1}|\gamma)$ along with a $\gamma$-invariant non-uniform choice of $C$ over $0,1]^{n}$ produces a joint distribution $F$ that is not conditionally independent but satisfies invariant dependencies. In terms of percentiles, the marginals $F^{j}$ governs how percentiles in each dimension are mapped into valuations in each dimension while $C$ governs the dependency structure between percentiles. As long as the dependencies between percentiles is invariant as $\gamma$ changes, any $\gamma$-dependent mapping from percentiles to valuations for each dimension satisfies invariant dependencies.
If $F$ has invariant dependencies then the $\mathbf{V}$ identified in (3) satisfies condition (1) and only the marginals matter. Given regularity along each marginal, pointwise maximization good-by-good solves the problem and produces the same optimal mechanism regardless of the dependency structure between goods as long as that dependency is invariant in $\gamma$.
Theorem 1 (Invariant Dependencies Implies Separability).
If $F$ has invariant dependencies, then the optimal joint mechanism is to offer the optimal sequential screening mechanism for each good separately.
@@
In the terminology of segal2014, $V$ is a vector of impulse responses of $\theta$ to $\gamma$. Recalling that $\nabla_{\theta}u(\gamma,\theta)=q(\gamma,\theta)$ gives the problem to be
$\mathbb{E}_{\gamma,\theta}\left[\sum_{j}q^{j}(\gamma,\theta)\phi^{j}(\gamma,\theta^{j})\right]\text{ for }\phi^{j}(\gamma,\theta^{j})=\theta^{j}+\frac{F_{\gamma}^{j}(\theta^{j}|\gamma)}{f(\theta^{j}|\gamma)}\frac{1-G(\gamma)}{g(\gamma)}.$
Here, $\phi^{j}$ coincides with the virtual value for good $j$ in a one-good sequential screening setting. When regularity holds, pointwise maximization produces a feasible solution; this pointwise maximization can be done good-by-good, leading to Theorem 1. Good-by-good the optimal mechanism consists of an upfront payment and an option value to buy; higher $\gamma$ consumers pay more upfront but have lower strike prices. Aggregating the optimal mechanism good-by-good leads to the optimal mechanism in a multidimensional screening setting to be one where the agent once again faces a menu of upfront payments (equal to the sum of upfront payments coming from the optimal mechanism for each good) and a separate strike price for each good.

### v2

@@ 4 Dynamic Decoupling
st. $\gamma\in\operatorname*{\arg\!\max}_{\gamma^{\prime}}\mathbb{E}_{\theta|\gamma}\left[u(\gamma^{\prime},\theta)\right]-t_{1}(\gamma^{\prime})$ $IC1$
@@
where $u(\gamma,\theta)=\theta\cdot q(\gamma,\theta)-t_{2}(\gamma,\theta)$ is the buyer’s utility net of $t_{1}$ and $\nabla_{\theta}u(\gamma,\theta)$ is the gradient of $u(\gamma,\theta)$ in $\theta$ at some fixed $\gamma$. Furthermore, a necessary condition for IC1 to hold is
$t_{1}(\gamma)=\int_{\Theta}u(\gamma,\theta)dF(\theta|\gamma)-\int_{\underline{\gamma}}^{\gamma}\int_{\Theta}u(\gamma^{\prime},\theta)f_{\gamma}(\theta|\gamma^{\prime})d\theta d\gamma^{\prime}$
which is used to arrive at the objective function.
@@
$IC1$ continues to pin down incentive compatibility for truthful reporting of $\gamma$ while $IC2$ captures incentive compatibility for truthful reporting of $\theta$. Next, $F$ is feasibility: The gradient of the buyer’s post-contractual utility function $u$ is equal to the allocation rule which must lie between zero and one. Finally, $IR$ pins down interim individual rationality to accept the contract. Total surplus generated by the monopolist is $\mathbb{E}_{\gamma,\theta}\left[\theta\cdot\nabla_{\theta}u(\gamma,\theta)\right]$ while a share $\mathbb{E}_{\gamma,\theta}\left[u(\gamma,\theta)\frac{f_{\gamma}(\theta|\gamma)}{f(\theta|\gamma)}\frac{1-G(\gamma)}{g(\gamma)}\right]$ is given to the buyer as information rents. The Lemma establishes that the interim utility function $u$ pins down both period one and period two transfers and specifies how dynamics distorts monopolist profits.
@@
Compared to the usual multi-good monopolist problem where the buyer receives their full information rents $u$, allowing the principal to frontload surplus extraction leads to period two information rents given to the buyer being extracted by period one transfers. As such, the only information rents the principal needs to pay the buyer are those to truthfully elicit their pre-contractual private information. These information rents are distorted by two factors. First, $\frac{f_{\gamma}(\theta|\gamma)}{f(\theta|\gamma)}=\frac{\partial}{\partial\gamma}\ln(f(\theta|\gamma))=s(\gamma,\theta)$ is the score function measuring how much the log-likelihood of $\theta$ changes as $\gamma$ changes. This captures how likely $\theta$ came from one value of $\gamma$ versus another and thus governs how post-contractual utilities are passed through into information rents. The other $\frac{1-G(\gamma)}{g(\gamma)}$ term is the standard hazard rate; as usual, providing information rents to lower types once again leads to higher information rents to higher types.
@@
The usual approach to multidimensional screening problems would be to apply the divergence theorem as in Daskalakis et al. (2017b), expressing $\nabla_{\theta}u(\gamma,\theta)$ in terms of $u(\gamma,\theta)$, and then applying duality. [^5] There are two complications preventing us from using an optimal transport or duality (see Kleiner and Manelli (2019)) approach to analyze the problem. First, as Pavan et al. (2014) notes, the consumer’s expected interim utility function may be discontinuous in $\gamma$. Unlike continuity in $\theta$ being guaranteed by feasibility of the allocation rule, the principal may choose to give information rents to consumers with different values of $\gamma$ at different values of $\theta$. While is is possible to use an orthogonalization approach to transform the problem, making $\mathbb{E}_{\theta|\gamma}\left[u(\gamma,\theta)\right]$ continuous in $\gamma$, it is unclear if $u(\gamma,\theta)$ itself can be made continuous in $\gamma$; or if doing so would be an useful exercise. Second, the usual single-crossing conditions that lead to monotonicity being a necessary and sufficient condition for implementability do not hold. We can view the problem as the principal designing a utility function $w$ where a buyer with private information $\gamma$ values $w$ at $\hat{U}(\gamma,w(\theta))=\int_{\Theta}w(\theta)dF(\theta|\gamma).$ Then, the usual characterizations of single crossing or monotonicity in $w$ cannot be applied as $w$ is a high-dimensional object. However, cyclic monotonicity as in Rochet (1987) is still a necessary and sufficient condition for implementability, but is difficult to operationalize. However, due to the additional distortion of $f_{\gamma}(\theta|\gamma)$ stemming from only paying the buyer information rents scaling with the informativeness of their initial private information, it is possible to instead apply the divergence theory to express $u(\gamma,\theta)$ in terms of $\nabla_{\theta}u(\gamma,\theta)$ in a simple way, especially so if invariant dependencies holds.
@@
$\nabla_{\theta}\cdot(\mathbf{V}(\theta|\gamma)f(\theta|\gamma))=f_{\gamma}(\theta|\gamma).$
@@
If invariant dependencies holds, $\mathbf{V}(\theta|\gamma)$ has a simple form: The $j$th component of $\mathbf{V}$ is
$\mathbf{V}^{j}(\theta|\gamma)=\frac{F^{j}_{\gamma}(\theta^{j}|\gamma)}{f^{j}(\theta^{j}|\gamma)}$
@@
and boundary conditions vanish. In the terminology of Pavan et al. (2014), $V$ is a vector of impulse responses of $\theta$ to $\gamma$. Recalling that $\nabla_{\theta}u(\gamma,\theta)=q(\gamma,\theta)$ gives the following formulation:
Lemma 2.
Suppose A3: Invariant Dependencies holds. Then, the monopolist’s problem can be written as:
$\max_{q(\gamma,\theta)}\text{ }$ $\mathbb{E}_{\gamma,\theta}\left[\sum_{j}q^{j}(\gamma,\theta)\phi^{j}(\gamma,\theta^{j})\right]$
st. $q(\gamma,\theta)\text{ is cyclically monotone in }\theta;$
$q(\gamma,\theta)\in[0,1]^{n}\text{ for all }\gamma,\theta;$
$IC1,IR.$
where
$\phi^{j}(\gamma,\theta^{j})=\left(\theta^{j}+\frac{F_{\gamma}^{j}(\theta^{j}|\gamma)}{f^{j}(\theta^{j}|\gamma)}\frac{1-G(\gamma)}{g(\gamma)}\right)$
denotes virtual value for good $j$.
If $F$ has invariant dependencies then only the marginals matter for the objective function. Then, taking the standard approach and ignoring global incentive compatibility constraints, maximizing the objective pointwise (i.e. allocating the good whenever virtual value is positive) leads to an allocation rule $q(\gamma,\theta)$ for which the allocation rule for good $j$ does not depend on the valuation for any other good $k\neq j$. When, in addition to A3, assumptions A1 and A2 also hold, the solution to the relaxed problem also satisfies global incentive constraints: Assumption A2 gives that the optimal allocation rule is increasing in $\gamma$ (and $\theta$) while Assumption A1 implies that buyer interim utility has increasing differences in $\gamma,q$ so any increasing allocation rule satisfies global incentive constraints. Since the relaxed problem in dimension $j$ exactly coincides with the relaxed problem if only good $j$ were present, the optimal mechanism simply is to run separate sequential screening mechanisms for each good.
Theorem 1 (Optimality of Separate Sales).
Under Assumptions A1-A3, the optimal joint mechanism is to offer the optimal sequential screening mechanism for each good separately: A buyer reporting $\gamma$ receives a price for good $j\in\{1,...,n\}$ of
$p^{j}(\gamma)=\min\left\{p:\phi^{j}(\gamma,p)\geq 0\right\}$
and pays an upfront cost of
$t_{1}(\gamma)=\sum_{j=1}^{n}\left[\int_{p^{j}(\gamma)}^{\bar{\theta}^{j}}1-F^{j}(\theta^{j}|\gamma)d\theta^{j}+\int_{\underline{\gamma}}^{\gamma}\int_{p^{j}(\gamma^{\prime})}^{\bar{\theta}^{j}}F_{\gamma}^{j}(t|\gamma^{\prime})dtd\gamma^{\prime}\right].$
As a corollary, the sequential screening mechanism posited in the motivating example maximizes monopolist surplus. Recall the setting: There are two goods with valuations $\theta^{1}=v-\varepsilon,\theta^{2}=v+\varepsilon$ for $v$ uniform over $[0,1/2]$ and $\varepsilon$ uniform over [1/2,1/2][-1/2,1/2]. The value $v$ is the buyer’s pre-contractual private information and $\varepsilon$ is the buyer’s post-contractual private information. Thus, final valuations are both uniform over [v1/2,v+1/2][v-1/2,v+1/2] and perfectly anti-correlated. In the language of the model, this corresponds to $v=\gamma\sim U[0,1/2]$ and $F^{1}(\theta^{1}|\gamma)=F^{2}(\theta^{2}|\gamma)=\max\{1-(\gamma+1/2-\theta),0\}$. As such, virtual values in this setting are
$\phi^{i}(\theta^{i}|\gamma)=\theta^{i}+\left(-\frac{1-2\gamma}{2}\right)=\theta^{i}-(1/2-v)$
so selling whenever virtual value is non-negative corresponds to posting prices of 1/2v1/2-v for each good. Period one transfers can then be computed to be $2v^{2}$. The allocation rule can be plotted as follows:
$\varepsilon\quad 0$1/21/21/2-1/201/41/41/2v1/2\ v(0,0)(0,0)(0,1)(0,1)(1,0)(1,0)$(1,1)$
4.1 Relaxing A3
To illustrate the complications that arise without invariant dependencies, consider the illustrative example with the following change: Let valuations remain perfectly anticorrelated for $\gamma<1/4$ but have valuations be perfectly correlated for $\gamma\geq 1/4$:
$\theta^{1}=v+\varepsilon,\theta^{2}=v-\varepsilon\text{ if }v<1/4$
$\theta^{1}=v+\varepsilon,\theta^{2}=v+\varepsilon\text{ if }v\geq 1/4$
where
$v\sim U[0,1/2];\varepsilon\sim U[-1/2,1/2].$
The previous mechanism of charging $2v^{2}$ upfront to secure strike prices of 1/2v1/2-v is still implementable and yields the following allocation rule:
$\varepsilon\quad 0$1/21/21/2-1/201/41/41/2v1/2\ v(0,1)(0,1)(0,0)(0,0)(1,0)(1,0)(1,1)(1,1)$(0,0)$
Note that if $v<1/4$ no consumer buys the grand bundle while if $v\geq 1/4$ consumers only buy the grand bundle. Thus, while strike prices of 1/2v1/2-v implements this allocation, the monopolist can do better via the following mechanism, charging a premia for the bundle:
  1. 1.
    If $v<1/4$: Pay $2v^{2}$ upfront for the right to purchase one (but not both) goods at a price of 1/2v1/2-v.
  2. 2.
    If $v\geq 1/4$: Pay $2v^{2}+1/8$ upfront for the right to purchase the bundle of goods at a price of 12v1-2v.
The only new incentive constraint introduced is along the border of $v=1/4$. Clearly, there will not be any upward deviations as a buyer with $v=1/4-\varepsilon$ cannot make any surplus from buying their less-preferred good at a price of 1/21/4=1/41/2-1/4=1/4. A buyer with $v=1/4$ receives surplus of
$\underbrace{2\cdot\frac{1}{2}\cdot(1/4+1/4)^{2}}_{\text{expected downstream gains}}-\underbrace{2(1/4)^{2}}_{\text{upfront payment}}-\underbrace{1/8}_{\text{premia}}=1/8-1/8=0$
from being truthful. If they report $1/4-\varepsilon$, their profit, up to $\varepsilon$ terms, is:
$\underbrace{\frac{1}{2}(1/4+1/4)^{2}}_{\text{single good}}-\underbrace{2(1/4)^{2}}_{\text{upfront payment}}=0.$
As such, this alternative mechanism satisfies incentive and individual rationality constraints. The monopolist is able to charge an extra $1/8$ with probability $1/2$ (i.e. whenever $v>1/4$) and their profit is thus (recall that profit from the original mechanism was $1/3$):
$\frac{1}{3}+\frac{1}{2}\cdot\frac{1}{8}=\frac{19}{48}.$
Total surplus is unchanged at $1/2$ so consumer surplus is decreased (by exactly the $1/16$ that the monopolist can now additionally extract):
$\frac{1}{2}-\frac{19}{48}=\frac{5}{48}<\frac{8}{48}=\frac{1}{6}.$
A very rough intuition is as follows: Since the correlational structure changes, the set of options in the monopolist’s menu also changes. This allows the monopolist to “reset” information rents at $v=1/4$. Buyers with $v=0$ and $v=1/4$ both receive zero surplus in expectation.
What is the optimal mechanism? To illustrate the forces involved, let us return to the general setting for now. Suppose $\Gamma$ can be partitioned into
$\Gamma=[\underline{\gamma},\overline{\gamma}]=[\underline{\gamma},\gamma_{1}]\cup[\gamma_{1},\overline{\gamma}]$
such that invariant dependencies holds over $\underline{\gamma},\gamma_{1}]$ and $\cup[\gamma_{1},\overline{\gamma}]$ separately. In an abuse of notation, the mechanism will be doubly-defined for $\gamma=\gamma_{1}$ since it is in both partitions; this is measure zero and is done to avoid infinitesimal terms. In this setting, the problem can be written as follows: The monopolist seeks to maximize
$\int_{[\underline{\gamma},\gamma_{1}]}\left[\int_{\Theta}\theta\cdot\nabla_{\theta}u(\gamma,\theta)d\theta-\int_{\underline{\gamma}}^{\gamma}\int_{\Theta}u(\gamma^{\prime},\theta)f_{\gamma}(\theta|\gamma^{\prime})d\theta d\gamma^{\prime}\right]g(\gamma)d(\gamma)$
$+\int_{[\gamma_{1},\overline{\gamma}]}\left[\int_{\Theta}\theta\cdot\nabla_{\theta}u(\gamma,\theta)d\theta-U_{1}-\int_{\gamma_{1}}^{\gamma}\int_{\Theta}u(\gamma^{\prime},\theta)f_{\gamma}(\theta|\gamma^{\prime})d\theta d\gamma^{\prime}\right]g(\gamma)d(\gamma)$
where $u(\gamma,\theta)$ is the utility function the monopolist designs and $U_{1}$ is the utility of an agent with $\gamma=\gamma_{1}$ at the threshold. For buyers with $\gamma\in[\underline{\gamma},\gamma_{1}]$, this term is the same as before. However, for buyers with $\gamma\in[\gamma_{1},\overline{\gamma}]$, the Fundamental Theorem of Calculus is applied with $\gamma_{1}$ at the lowest point instead of $\underline{\gamma}$; this leads to baseline utility being $U_{1}$ instead of 0 but the integral starting at $\gamma_{1}$ instead of $\underline{\gamma}$. The new constraints are incentive compatibility and individual rationality at the boundary ($U_{1}\geq 0$). Applying the tools developed in this paper to each interval separately gives the following objective:
$\mathbb{E}_{\gamma\in[\underline{\gamma},\gamma_{1}],\theta}\left[\sum_{j}q^{j}(\gamma,\theta)\left(\theta^{j}+\frac{F_{\gamma}(\theta^{j}|\gamma)}{f^{j}(\theta^{j}|\gamma)}\frac{G(\gamma_{1})-G(\gamma)}{g(\gamma)}\right)\right]$
$+\mathbb{E}_{\gamma\in[\gamma_{1},\overline{\gamma}],\theta}\left[\sum_{j}q^{j}(\gamma,\theta)\left(\theta^{j}+\frac{F_{\gamma}(\theta^{j}|\gamma)}{f^{j}(\theta^{j}|\gamma)}\frac{1-G(\gamma)}{g(\gamma)}\right)\right]-U_{1}[1-G(\gamma_{1})].$
There are two forces that go in opposite directions. The standard “no distortions at the top” intuition manifests itself in allocating efficiently to the highest type in $\underline{\gamma},\gamma_{1}]$ as their virtual value for good $j$ is
$\theta^{j}+\frac{F_{\gamma}(\theta^{j}|\gamma)}{f^{j}(\theta^{j}|\gamma)}\frac{G(\gamma_{1})-G(\gamma=\gamma_{1})}{g(\gamma)}=\theta^{j}.$
This suggests that the monopolist should allocate more of the good than before to buyers in $\underline{\gamma},\gamma_{1}]$. On the other hand, if too much of the good is allocated to agents at the high end of $\underline{\gamma},\gamma_{1}]$, the value of $U_{1}$ must increase to deter downward deviations. This alternative force suggests that the monopolist should allocate less of the good than before to buyers in $\underline{\gamma},\gamma_{1}]$. How these forces are balanced also depends on how the distribution of $\theta$ discontinuously changes at $\gamma_{1}$. Furthermore, there are multiple mechanisms where both individual rationality and incentive compatibility at the border are binding; the difficulties associated with multidimensional screening come back in this case. Another takeaway is that the optimal mechanism for the upper partition remains unchanged.
Return to the illustrative example with these intuitions: The monopolist would like to allocate more of the good to buyers in $[0,1/4]$ while preserving giving the lowest type in $[1/4,0]$ zero information rents. Using $G(1/4)$ instead of $G(1/2)=1$ in the expression for virtual values for buyers in $[0,1/4]$ gives their virtual value to be
$\theta^{i}+\left(\frac{-1/2-2\gamma}{2}\right)=\theta^{i}-(1/4-v)$
which suggests strike prices of 1/4v1/4-v. However, buyers in $[1/4,1/2]$ value the option of purchasing a bundle much more than buyers in $[0,1/4]$ since their values are perfectly correlated (instead of being perfectly anti-correlated). This suggests the following mechanism (period one transfers are calculated using the standard formula for transfers):
  1. 1.
    If $0\leq v\leq 1/8$: Pay $2v^{2}+\frac{1}{2}v+\frac{1}{16}$ to secure a strike price of 1/4v1/4-v for one good.
  2. 2.
    If $1/8\leq v\leq 1/4$: Pay $\frac{9}{32}$ upfront to claim exactly one of the goods for free.
  3. 3.
    If $1/4\leq v$: Pay $2v^{2}+1/8$ upfront for the right to purchase the bundle of goods at a price of 12v1-2v.
Case one simply reflects the pointwise optimal mechanism. Case two is the optimal mechanism subject to not selling a bundle. Note that at $v=1/8$, strike prices are 1/41/8=1/81/4-1/8=1/8 so the consumer will always purchase one of the goods. As such, it is without loss to charge them the $1/8$ upfront; this is added to the standard upfront fee to arrive at $9/32$. Charging them upfront instead of downstream helps with preventing downward deviations from buyers in $[1/4,1/2]$: These buyers will not purchase either good at a price of $1/8$ with some probability so frontloading the transfer hurts their deviation payoff. Finally, case three is the optimal mechanism from before with a bundle premia to set the utility of the lowest type in $[1/4,1/2]$ to be equal zero. This induces the following allocation rule:
$\varepsilon\quad 0$1/21/21/2-1/201/41/41/2v1/2\ v1/81/8(0,0)(0,0)(0,1)(0,1)(1,0)(1,0)(1,1)(1,1)$(0,0)$
Finally, we verify incentive constraints around $1/4$. A buyer in $[0,1/4]$ deviating upwards will pay $2(1/4)^{2}+1/8=1/4$ to have the option to pay 12(1/4)=1/21-2(1/4)=1/2 to receive a bundle they value at $(1/4-\varepsilon)+(1/4+\varepsilon)=1/4$. Thus their expected payoff is 1/4-1/4 but their payoff from being truthful is strictly positive. A buyer in $[1/4,1/2]$ deviating downward will pay $9/32$ to get an item. The expected value of an item is
$\mathbb{E}_{\varepsilon\sim U[-1/2,1/2]}\left[\max{1/4+\varepsilon,0}\right]=9/32$
so their deviation payoff is $9/32$. Since $q$ is increasing in $\gamma$ separately over $[0,1/4]$ and $[1/4,1/2]$, larger deviations between $[0,1/4]$ and $[1/4,1/2]$ can be ruled out as well. As such, this mechanism is incentive compatible. Monopolist profits are
$\frac{157}{384}>\frac{152}{384}=\frac{19}{48}$
while total surplus is
$\frac{71}{128}>\frac{64}{128}=\frac{1}{2}$
and consumer surplus is
$\frac{7}{48}>\frac{5}{48}.$
As such, this mechanism, which further exploits the change in correlational structure by forcing low-value buyers to only acquire a single good, increases both firm and consumer surplus.
It remains an open question if this is the globally optimal mechanism. Both individual rationality and incentive compatibility constraints bind at the boundary of $v=1/4$. It generally is impossible to give individuals in $[0,1/4]$ more of the good without keeping information rents of the lowest type in $[1/4,1/2]$ at zero.
4.2 Relaxing A2 or A1
Assumption A2: Regularity establishes that the pointwise maximizer of the relaxed problem is an allocation rule that is increasing in both $\gamma$ and $\theta^{j}$. Allocation rules being increasing in $\theta^{j}$ is necessary and sufficient (given additive separable sales) for period two incentive constraints to hold. Assumption A1: Dimensions are commonly first-order stochastic dominance-ordered implies that any allocation rule being weakly increasing in $\gamma$ satisfies global incentive constraints.
There are two dimensions to relax this pair of constraints: less regularity in $\gamma$ or less regularity in $\theta$. Without regularity in $\gamma$, different goods become endogenously linked via which good(s) are utilized to provide period one incentives. For example, suppose that for some subset of goods, either (1) higher values of $\gamma$ led to FOSD-lower distributions of values or (2) the pointwise solution to the relaxed problem was decreasing in $\gamma$. Either of these effects would provide an incentive for the buyer to deviate to collect higher surplus from transacting on that subset of goods. However, the usual argument applied to other goods implies that the buyer would hurt otherwise. If violations of A1 or A2 were sufficiently small on some subset of goods, the pointwise solution remains implementable. Otherwise, the monopolist may need to balance which dimensions are used to provide incentives.
What about regularity in $\theta$? A natural question would be when ironing separately in each dimension is sufficient.
Proposition 1 (Separate Ironing).
Suppose $\theta^{i}$ is independent of $\theta^{j}$ for all $i,j$ and
$\theta^{j}f^{j}(\theta^{j}|\gamma)g(\gamma)+F_{\gamma}(\theta^{j}|\gamma)G(\gamma)$
is increasing in $\gamma$. Then, ironing each dimension independently produces the optimal mechanism.
Independence is necessary for the problem to be strongly decomposable in each dimension. Without independence, a cyclically monotone $q(\gamma,\theta)$ may induce interim allocations
$\hat{q}^{j}(\gamma,\theta^{j})=\mathbb{E}_{\theta^{-j}}\left[q(\gamma,\theta)\right]$
that are not monotone in $\theta^{j}$ based on how valuations are coupled; this holds even if assumption A3 is in effect. With independence, every cyclically monotone $q$ induces monotone allocation rules in each dimension so it is without loss to set $q^{j}$ solely as a function of $\gamma,\theta^{j}$. The other assumption supposes regularity on virtual values weighted by the probability of realizing $(\gamma,\theta)$ pairs. Under this condition, the optimization problem in each dimension has increasing differences in $q,\gamma$ so higher values of $\gamma$ yield pointwise-higher allocation rules, regardless of how ironing is done.
5 Extensions
To test the limits of how powerful frontloading surplus extraction is, we compare the model studied in this paper to two other benchmarks. First, one additional period is all the monopolist needs: If values are independent and private information for each good comes over time (as in Pavan et al. (2014)), the monopolist cannot utilize future allocations to extract additional surplus. Here, we need independence and not just invariant dependencies since if information comes over time, any dependencies can give the monopolist more power, thus stacking the deck too much in the monopolist’s favor. Second, we suppose that the monopolist is able to directly observe an orthogonalized signal of the buyer’s post contractual information (as in Eső and Szentes (2017)). Once again, the monopolist cannot benefit, providing another interpretation of frontloading surplus extraction allowing the monopolist to avoid paying any post-contractual information rents. Third, we consider the case where there are $n$ items, each in unit supply, to be auctioned as in Eső and Szentes (2007). In this case, the optimal auction for the $n$ items is to host separate sequential screening auctions for each good.